### v1

# AKRMap: Adaptive Kernel Regression for Trustworthy Visualization of Cross-Modal Embeddings

Abstract
Cross-modal embeddings form the foundation for multi-modal models. However, visualization methods for interpreting cross-modal embeddings have been primarily confined to traditional dimensionality reduction (DR) techniques like PCA and t-SNE. These DR methods primarily focus on feature distributions within a single modality, whilst failing to incorporate metrics (e.g., CLIPScore) across multiple modalities. This paper introduces AKRMap, a new DR technique designed to visualize cross-modal embeddings metric with enhanced accuracy by learning kernel regression of the metric landscape in the projection space. Specifically, AKRMap constructs a supervised projection network guided by a post-projection kernel regression loss, and employs adaptive generalized kernels that can be jointly optimized with the projection. This approach enables AKRMap to efficiently generate visualizations that capture complex metric distributions, while also supporting interactive features such as zoom and overlay for deeper exploration. Quantitative experiments demonstrate that AKRMap outperforms existing DR methods in generating more accurate and trustworthy visualizations. We further showcase the effectiveness of AKRMap in visualizing and comparing cross-modal embeddings for text-to-image models. Code and demo are available at https://github.com/yilinye/AKRMap.
Machine Learning, ICML
1 Introduction
Figure 1: CLIPScore distribution on the COCO dataset by t-SNE (a). The visualization shows dense neighboring points with significantly different metric values, causing overlapping and occlusion (b) and highly inaccurate contour mapping (c).
Cross-modal embeddings play a fundamental role for multi-modal models, functioning as cross-modal encoders (Rombach et al., 2022), objective functions (Ramesh et al., 2022), or evaluation metrics (Hessel et al., 2021; Wu et al., 2023a) for tasks like text-to-image (T2I) generation. To assess the alignment of multi-modal models, various evaluation metrics based on these embeddings have been introduced, such as CLIPScore (Hessel et al., 2021), and Human Preference Score (HPS) (Wu et al., 2023a, b; Zhang et al., 2024). Despite their utility, embedding-based evaluations are often difficult to interpret, as the metrics are typically reported as aggregated values, without providing insights into instance-level performance. This deficiency undermines the trustworthiness and transparency of the evaluation.
To address this limitation, dimensionality reduction (DR) visualization serves as a crucial tool, offering a way to reveal the landscape of cross-modal metrics and enabling a more comprehensive understanding of model performance. Recent studies (Liang et al., 2022; Wang et al., 2023b, c) have explored the use of DR methods for cross-modal embeddings. These efforts often leverage established techniques such as PCA (Li et al., 2018), UMAP (McInnes et al., 2018) and t-SNE (Van der Maaten & Hinton, 2008), as well as autoencoder-based approaches (Le et al., 2018; Elhamod & Karpatne, 2024). However, existing DR methods are primarily designed to depict feature distributions within a single modality. When applied to cross-modal metrics, these methods often produce dense neighborhoods where points with significantly different metric values are positioned close, leading to overlaps and local occlusion (Figure 1(b)). Moreover, multi-modal models are typically evaluated on large-scale datasets containing millions of data points. This creates a need for rendering contour maps that can reveal continuous metric distributions. However, local neighboring points with a mix of high and low metric values may be misrepresented, as the contour map depicts only a single aggregated value, causing inaccurate contour mapping as illustrated in Figure 1(c).
To construct trustworthy visualizations for cross-modal embeddings, a key consideration is to enhance the accuracy of metric contour mapping in the projected 2D space. As contour estimation typically relies on radius basis function (RBF) kernel (e.g., Gaussian kernel), we seek to enhance the kernel-based mapping by coordinating it with the DR process. Drawing inspiration from supervised t-SNE (Gisbrecht et al., 2015) and UMAP (Sainburg et al., 2021), we introduce Adaptive Kernel Regression Map (AKRMap), a supervised DR method that leverages adaptive kernel regression to effectively visualize the distribution of cross-modal embedding metrics. Specifically, AKRMap first constructs a supervised DR network explicitly guided by post-projection kernel regression loss with neighborhood constraint (Sect. 4.1.1). Next, to account for the gap between high-dimensional and low dimensional kernels, we improve the flexibility of the post-projection kernel through adaptive generalized kernel jointly learned with the projection (Sect. 4.1.2). This approach enables the generation of scatterplots for discrete data points and contour maps for continuous metric landscapes, while also supporting advanced features such as overlay views and zooming for interactive interpretation (Sect. 4.2). Both the DR method and the visualization tool have been implemented as a Python package and will be made publicly available.
We conduct quantitative experiments to evaluate AKRMap against traditional and autoencoder-based DR visualizations (Sect. 5.1). The results highlight the superior performance of AKRMap in accurately mapping cross-modal metrics for both in-sample and out-of-sample data points. We further demonstrate the practical applications of AKRMap across three distinct scenarios (Sect. 5.2): 1) visual exploration of human preference dataset (HPD) (Wu et al., 2023a), 2) visual comparison of diffusion-based and auto-regressive T2I models, and 3) visual examination of the global impact for fine-tuning. These applications illustrate that AKRMap effectively enables human-in-the-loop interpretation of model performance on large-scale datasets.
In summary, our contributions are as follows:
  • We propose a novel DR method for cross-modal metric visualization, which jointly learns the projection and metric contour mapping through kernel regression supervised DR with adaptive generalized kernel.
  • We develop a tool for trustworthy visualization of cross-modal metrics, incorporating visualization features such as a scatterplot view and contour maps, along with interactive features like zooming and overlaying.
  • We conduct quantitative experiments to demonstrate the superior performance of AKRMap in generating more accurate visualizations of cross-modal metric, and highlight its applications across three scenarios to enhance the trustworthiness of T2I model evaluations.
2 Related Work
2.1 Cross-Modal Embedding-based Evaluation
Cross-modal embeddings like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) form the foundation of multi-modal learning. Specifically, many evaluation methods for multi-modal models, such as T2I models, rely on these embeddings. Conventional metrics like Inception Score (Salimans et al., 2016) and Fréchet inception distance (Heusel et al., 2017) compute the average distance or distributional difference between generated and reference images in the embedding space, yet they are insufficient to measure cross-modal alignment and instance-level performance. With the advancement of multi-modal AI, cross-modal embedding metrics like CLIPScore (Hessel et al., 2021) have emerged to measure the alignment between prompts and generated images. While being effective in capturing the alignment, CLIPScore fails to model human preferences. To address this limitation, recent studies have proposed training specialized human preference models such as HPS (Wu et al., 2023a, b; Zhang et al., 2024), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2024), to better align with human judgments. For example, ImageReward (Xu et al., 2024) introduces a systematic pipeline for expert annotations to train preference models, while PickScore (Kirstain et al., 2023) leverages large-scale crowd-sourced human comparisons.
However, these metrics generally provide only an average score, offering a broad overview of a model’s performance across the entire embedding space. This highlights the need for neural embedding visualizations that can expose the detailed distribution of metric values within the cross-modal embeddings, facilitating instance-level inspection and supporting human-in-the-loop evaluation.
Figure 2: AKRMap is a neural network based DR method designed to learn adaptive kernel regression for visualizing cross-modal embeddings. The network integrates two key components to jointly learn data point projections and cross-modal metric estimation: 1) Kernel regression supervision, and 2) Adaptive generalized kernel. The resulting visualizations, including scatterplots and contour maps, provide a clearer and more accurate representation of the cross-modal metric distribution.
2.2 Visualization for Neural Embeddings
Visualization is an important tool to support data analysis in conjunction with AI (Wang et al., 2023a; Ye et al., 2024). Particularly, it has proven to be effective for enhancing the understanding of various types of neural embeddings, including word embeddings (Mikolov et al., 2013; Chiang et al., 2020), vision-language embeddings (Liang et al., 2022; Wang et al., 2023b; Ye et al., 2025), and parameter spaces within loss landscape (Li et al., 2018; Elhamod & Karpatne, 2024). Commonly, dimensionality reduction (DR) techniques are employed for visualization. These include linear DR methods such as PCA (Abdi & Williams, 2010), as well as nonlinear DR methods like t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018). Parametric versions of traditional DR methods, which provide explicit mapping for projections, can be achieved by training neural networks, such as parametric t-SNE (Gisbrecht et al., 2015; Damrich et al., 2023), parametric UMAP (Sainburg et al., 2021) and other parametric methods for interactive clustering (Xia et al., 2023) or streaming data (Xia et al., 2024). Similarly, auto-encoder based DR methods (Le et al., 2018; Elhamod & Karpatne, 2024) have also been developed, incorporating an additional objective of reconstructing high-dimensional embeddings from their projections. Recently, there has been a growing interest in utilizing DR techniques to visualize cross-modal embeddings. For instance, UMAP has been employed to explore and visualize the modality gap between text and image embeddings (Liang et al., 2022; Wang et al., 2023b). To address this modality gap and enhance the cross-modal alignment in visualization space, some studies have introduced fusion-based DR methods (Ye et al., 2025), enabling the visualization of image embeddings in relation to text embedding anchor points.
However, previous efforts treat multi-modal embeddings separately, failing to effectively depict the landscape of cross-modal metrics (e.g., CLIPScore and HPS). In addition, existing metric-aware DR techniques focus on a limited range of particular metrics like distance (Sainburg et al., 2021) and density (Narayan et al., 2021), without a generalizable mechanism to preserve other different metrics. To address this limitation, we propose a kernel regression-based supervised projection method combined with an adaptive generalized kernel. Different from traditional kernel DR methods like kernel PCA (Schölkopf et al., 1997) that mainly use kernel to transform the high-dimensional features, we develop an adaptive kernel in the projection space to guide the DR. Notably, to construct the kernel guidance, we propose a novel cross-validation supervision trick that bridges between traditionally nonparametric kernel regression and parametric DR. This design allows the propagation of contour estimation errors back into the projection process. In this way, our method can dynamically adjust both the DR mapping and the kernel shape, enabling it to accurately model and fit the complex landscape of cross-modal metrics.
3 Problem Definition
Taking text-to-image generation models as an example, we formally define embedding-based metric for multi-modal models: A pretrained neural network $e(\cdot)$ encodes the prompt $t$ and the generated image $v$ into a high-dimensional embedding representations $e(t,v)$, which are subsequently used to predict a metric score $s=f(e)$. Here, $f(\cdot)$ represents the final operation applied to the embeddings, such as cosine distance in HPS (Wu et al., 2023a).
The problem of generating trustworthy visualizations for cross-modal embeddings can be defined as follows: given a dataset of prompt-image pairs $D=\{(t_{i},v_{i})\}_{i=1}^{N}$ and its embedding representations $e(D)=\{e(t_{i},v_{i})\}_{i=1}^{N}$, we first seek to learn a manifold $M$ (Goldberg et al., 2008) where $e(D)$ resides in the high-dimensional embedding space $R^{d}$. This manifold can be "spread out" (reparametrized) in the visualization space $R^{2}$ to show the distribution of metric score across the dataset, where the process can be modeled as a projection mapping $P(\cdot)\ s.t.\ P(e(D))\in R^{2}$. This process aims to learn an explicit mapping function that projects high-dimensional embeddings to points in 2D space while accurately preserving the underlying metric distribution.
Next, to visualize a continuous distribution from the discrete projected sample points in $P(e(D))$, a contour map in the 2D space needs to be estimated. Here, we adopt the approach used by contouring algorithms in various Python libraries, such as Plotly, by dividing the projected 2D space into a grid and calculating the metric distribution values at each grid point. Specifically, suppose the projected coordinates lie within a normalized 2D space of $[0,1]\times[0,1]$. For each grid $\mathbb{x}_{g}=(\frac{i}{N_{w}},\frac{j}{N_{h}})$, we compute a value $\hat{s}(\mathbb{x}_{g})$ based on the local metric distribution, which is then colored using a continuous colormap. $N_{w}$, $N_{h}$ denotes the number of grids along the x-axis and y-axis, respectively. In this manner, we can generate a contour map depicting continuous landscape of the metric distribution from the discrete projected sample points.
4 AKRMap
We propose Adaptive Kernel Regression Map (AKRMap), with the workflow illustrated in Figure 2. The input is a set of cross-modal embeddings with each high-dimensional vector representing a data point in the embedding space. First, to synchronize the contour mapping with the projection for accurate metric landscape estimation, we propose a cross-modal metric-supervised projection method comprising two key components: 1) Kernel Regression Supervision, which guides the projection to achieve more precise metric mapping (Sect. 4.1.1), and 2) an Adaptive Generalized Kernel, which accounts for the gap between high-dimensional and low-dimensional kernels and allows for more flexible contour mapping to capture complex metric landscapes (Sect. 4.1.2). Then, we design an interactive visualization tool to facilitate multi-scale exploration of metric distribution and individual data points (Sect. 4.2).
4.1 Adaptive Kernel Regression
4.1.1 Kernel Regression Supervised Projection
According to the problem definition, to achieve an accurate cross-modal metric visualization, cross-modal metric supervision is necessary. To address the challenge of constructing appropriate supervision for continuous cross-modal metric, we develop a kernel regression supervised projection method. Specifically, we propose learning a projection network $P:\ R^{n}\rightarrow R^{2}$ with a Nadaraya-Watson kernel regression (Ali et al., 2023) in the projected space:
$\hat{s}(\mathbb{x})=\frac{\sum_{k=1}^{N}K(\mathbf{x}-P(e(t_{k},v_{k})))\cdot s_{k}}{\sum_{k=1}^{N}K(\mathbf{x}-P(e(t_{k},v_{k})))},$ (1)
where $P(e(t_{k},v_{k})))$ is the projected sample point of metric embeddings and $s_{k}$ is the corresponding ground-truth metric value in the training set. $K(\cdot)$ is an RBF kernel in $R^{2}$.
Subsequently, to construct a supervised learning objective for the projection, inspired by the cross-validation method for kernel learning (Silverman, 2018), we randomly split the dataset $D$ into training set $D_{tr}$ and validation set $D_{vl}$ of ratio $9:1$ in each epoch. Then, in equation (1), we estimate the metric distribution only with points in training set ($(t_{k},v_{k})\in D_{tr}$). Next, we seek to minimize the weighted mean square error loss:
$MSE_{p}=\frac{1}{|D_{p}|}\sum_{x_{i}\in D_{p}}(\hat{s}(x_{i})-s_{i})^{2},\ p\in\{vl,tr\},$ (2)
$MSE_{r}=w_{1}MSE_{vl}+w_{2}MSE_{tr},$ (3)
where the overall regression loss $MSE_{r}$ is a weighted sum of the loss on validation set ($MSE_{vl}$) and that on training set ($MSE_{tr}$). We seek to balance these loss terms to ensure mapping accuracy at both in-sample and out-of-sample positions on the contour, and the weights $w_{1}=1$ and $w_{2}=0.3$ are set empirically.
Neighborhood Preservation Constraint. To encourage the projection to maintain some neighborhood information of the high-dimensional embeddings, we combine our regression loss with a constraint term from traditional dimensionality reduction. Specifically, we incorporate the KL-divergence loss from t-SNE:
$KL=\sum_{i}^{n}\sum_{j,j\neq i}^{n}p_{ij}\ln\frac{p_{ij}}{q_{ij}},$ (4)
where $p_{ij}$ and $q_{ij}$ are the neighborhood distribution probabilities in high-dimensional and low-dimensional space respectively, as detailed in (Van der Maaten & Hinton, 2008). Specifically, we adopt a perplexity-free implementation (De Bodt et al., 2018; Crecchi et al., 2020) of the KL loss for our parametric projection network. Overall, the objective function of our method is:
$L=\lambda MSE_{r}+KL,$ (5)
where $\lambda=0.125$ is empirically set to balance the loss.
4.1.2 Adaptive Generalized Kernel
In this section we illustrate the design of kernel $K(\cdot)$ in the regression supervision in equation (1). The key challenge is that we need to consider the discrepancy between high-dimensional and low-dimensional kernels. First, to avoid unstable optimization caused by exponentials and make the regression more compatible with our low-dimensional t-SNE neighborhood constraint, we adopt a t-distribution-like kernel instead of a Gaussian kernel. However, for complex metric distribution landscapes, especially those in embedding-based generative model evaluation, a standard kernel may lack enough flexibility. Particularly, it is difficult to determine a proper decay rate of the kernel value in relation to the distance. To address the problem, we take inspiration from the observation in a previous study (Narayan et al., 2021) of an approximate power law relationship between the local radius of the low-dimensional space ($r_{e}$) and that of original high-dimensional space $r_{o}$: $r_{e}\left(x_{i}\right)\approx\alpha\left[r_{o}\left(x_{i}^{h}\right)\right]^{\beta}$. We hypothesize that this transformation can also improve the adaptability of low-dimensional kernel. Therefore, we propose using an adaptive generalized t-distribution kernel:
$K(\mathbf{x},\alpha,\beta)=\left(1+\alpha\left\|\mathbf{x}\right\|^{2\beta}\right)^{-1},$ (6)
where $\alpha$ and $\beta$ are learnable parameters that are jointly optimized with the projection. In this way, our method can dynamically change the shape of the kernel to accurately fit the cross-modal metric landscape and reduce the risk of overfitting or underfitting due to suboptimal decay rate.
4.2 Cross-modal Metric Visualization
On the basis of AKRMap, we further develop an interactive visualization tool that provides two distinct views for exploring cross-modal embedding metrics:
  • Scatterplot view. The scatterplot view displays all individual data points within the embedding space, with each point color-coded according to the cross-modal embedding metric. This view allows for direct interaction with the data points. Unlike scatterplots generated by baseline DR methods, which often suffer from significant occlusion that obscures the true metric distribution, AKRMap effectively reveals extreme values with greater clarity, as demonstrated in Figure 3.
  • Contour map. For large datasets like HPD and COCO, scatterplot can become overcrowded with cluttered points. To address this limitation, we introduce a contour map that effectively represents cross-modal metric distribution in a continuous manner, by dividing the 2D space into grids and computing grid values based on the local metric distribution. This representation makes regions with extreme values more prominent and reveals distribution patterns with greater clarity.
Figure 3: Comparison of scatterplots generated by t-SNE and AKRMap for the HPSv2 metric. Despite the visual clutter introduced by the large-scale dataset, AKRMap provides a clearer and more accurate representation of the HPSv2 metric distribution.
The visualization tool includes various interactive features designed to facilitate in-depth exploration, including:
  • Zoom. AKRMap enables efficient multiscale zooming by dynamically computing the contour map using varying grid resolutions, as illustrated in Figure 4. Specifically, it adjusts the level of detail by increasing the grid resolution proportionally to the zoom level, revealing finer details in the map as users zoom in. This capability is unique to AKRMap and not achievable with traditional DR methods, which lack the ability to preserve and display local details.
  • Overlay. The scatterplot view and contour map allow data points to be overlaid onto the contour map, enabling users to simultaneously observe the overall distribution while interacting with individual data points, as shown in Figure 4. To mitigate occlusion, we have implemented sampling techniques, including random sampling and Poisson disk sampling.
The visualization tool has been built into an easy-to-use python package with minimal dependencies, requiring only PyTorch and Plotly, which can be seamlessly integrated into interactive computational notebooks.
Figure 4: Contour map combined with zoom and overlay with point sampling for multiscale exploration of the HPD dataset.
5 Experiment
Table 1: Quantitative comparison of cross-modal embedding metric visualizations on the HPD dataset. Our AKRMap outperforms baseline methods and ablations, achieving the best performance across all four cross-modal embedding metrics.
Method CLIPScore HPSv2 PickScore Aesthetic Score
$mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$
PCA 4.0042 16.5506 5.1142 1.4444 5.7870 1.8749 1.1497 5.9206 1.4686 0.5322 11.5645 0.6756
t-SNE 4.0241 16.6643 5.1523 1.4361 5.7568 1.8629 1.1579 5.9429 1.4640 0.4725 10.3128 0.6018
UMAP 4.0819 16.9066 5.2179 1.4432 5.7817 1.8725 1.2214 6.2721 1.5536 0.4463 9.6323 0.5721
SAE 4.1843 16.9386 5.2844 1.4483 5.8117 1.8696 1.2357 6.3952 1.5650 0.5809 12.7041 0.7251
Neuro-Visualizer 12.4201 43.8219 13.7318 1.6871 6.4526 2.0649 1.9655 9.5295 2.3427 0.4900 10.6306 0.6252
AKRMap (w/o KR) 4.0164 16.4830 5.1274 1.3978 5.6039 1.8102 1.1284 5.8004 1.4353 0.4752 10.3943 0.6003
AKRMap (w/o GK) 2.2812 9.1189 2.9458 1.1430 4.5498 1.4695 0.9064 4.6194 1.1686 0.4555 9.9267 0.5718
AKRMap 1.8707 7.3649 2.4253 0.8108 3.1935 1.1200 0.7712 3.9225 1.0225 0.4305 9.3340 0.5488
5.1 Quantitative Experiments
Baseline methods. We compare AKRMap against three commonly used DR methods: PCA, t-SNE and UMAP, as well as two encoder-based approaches: SAE (Le et al., 2018) and Neuro-Visualizer (Elhamod & Karpatne, 2024). For the encoder-based methods, the decoder is used to estimate grid colors, while the other baseline methods rely on the common Gaussian RBF kernel to estimate the 2D distribution map. We use distance threshold to cut off empty areas for traditional DR techniques but keep the full landscape of autoencoder methods because far away positions in traditional DR results are ambiguous and cannot hardly be mapped back to original high-dimensional space. Additional details are provided in Appendix C.
Ablation Study. We evaluate AKRMap under two alternative settings: 1) training the projection network using only the neighborhood constraint, without the supervision provided by Kernel Regression (AKRMap w/o KR), and 2) removing the adaptive Generalized Kernel component (AKRMap w/o GR).
Dataset. To evaluate performance on the cross-modal embedding metric, we select the widely used large-scale T2I dataset, the Human Preference Dataset (HPD) (Wu et al., 2023a). The HPD dataset contains over 430,000 unique images generated by various T2I models, along with their corresponding prompts, in the official training set, and 3,700 images in the test set.
Cross-modal Embedding Metrics. We compare AKRMap against baseline methods for visualizing several commonly used cross-modal embedding metrics in T2I generation, including CLIPScore (Hessel et al., 2021), HPSv2 (Wu et al., 2023a), PickScore (Kirstain et al., 2023), and Aesthetic Score (Schuhmann et al., 2022).
Performance Evaluation. We evaluate the accuracy and trustworthiness of the visualization methods using mapping errors, including $mae$, $rmse$, and $mape$. Specifically, we calculate these errors on the test set of HPD for out-of-sample points that were not used during training. We also report the errors for in-sample points ($mae_{in}$, $mape_{in}$, $rmse_{in}$) from the HPD training set in Appendix B. This dual evaluation is important because a good visualization method must accurately represent the training data distribution while avoiding misleading maps caused by overfitting and ensuring reliable accuracy for test points.
Training implementations. The architecture of our projection network is a 4-layer MLP, with shape $(d,d)$ in the first three layers and $(d,2)$ in the last layer, where $d$ is the dimensions of the input embeddings. Batch normalization and ReLU activation are applied to each layer. Our projection model is trained on one Nvidia L4 GPU with batch size of 1000 and 20 epochs. We use Adam optimizer with a learning rate of 0.002. For the t-SNE and PCA implementations, we use the python sklearn package, where the t-SNE method adopts the Barnes-Hut t-SNE (Van Der Maaten, 2014). For the UMAP implementation, we adopt the python umap-learn package. The Neuro-Visualizer implementation is based on the open-sourced code of the original paper (Elhamod & Karpatne, 2024), while the SAE is based on a github reimplementation [^1] https://github.com/mortezamg63/Supervised-autoencoder.
The experiment results are presented in Table 1.
Comparison with baselines. Our method AKRMap consistently outperforms baseline methods in mapping accuracy across all the embedding metrics. Notably, AKRMap effectively reduces mapping errors for both training (in-sample) points (see Appendix B) and test (out-of-sample) points, demonstrating its ability to produce more trustworthy mappings. For instance, when applied to the HPSv2 metric, AKRMap reduces the MAE by nearly 50% for in-sample points and approximately 43% for out-of-sample points compared to the best-performing baseline, t-SNE. In addition, while autoencoder methods are effective for loss landscape, it has shown weakness in cross-modal metric mapping with relatively higher error and unstable performance. These findings highlight AKRMap’s superior reliability and robustness across diverse embedding scenarios.
Ablation results. The ablation results highlight the importance of the two key components of our method: kernel regression supervision (KR) and the adaptive generalized kernel (GK), both of which are critical for enhancing mapping accuracy. Among these, kernel regression supervision proves to be the most impactful, as the results for the AKRMap w/o KR setting are nearly indistinguishable from those of the baseline methods in Table 1. This outcome is expected, as removing kernel regression supervision leaves the projection network relying solely on the neighborhood constraint, making it functionally similar to the t-SNE method. In addition, the adaptive generalized kernel demonstrates its value by further reducing errors across various metrics, underscoring its effectiveness in capturing and fitting complex metric landscapes. These findings validate the necessity of both components in achieving superior performance.
5.2 Applications
5.2.1 Visualizing Large T2I Dataset
Figure 5: Qualitative comparison of contour map visualizations of ClipScore, HPSv2, PickScore, and Aesthetic Score distributions in the HPD dataset generated by our AKRMap and four baselines: PCA, UMAP, t-SNE, and Neuro-Visualizer.
AKRMap can provide an accurate overview of large T2I dataset by effectively capturing the cross-modal metric distribution. Figure 5 shows the contour maps of ClipScore, HPSv2, PickScore, and Aesthetic Score distributions in the HPD dataset, generated by our AKRMap and four baseline methods: PCA, UMAP, t-SNE, and Neuro-Visualizer.
In alignment with the results from the quantitative experiments, our method demonstrates superior performance by accurately depicting score distributions with rich local details. In contrast, the baselines suffer from various limitations, such as over-smoothing due to averaging effects, which obscure local structures. For example, PCA introduces pronounced block effects with sharp edges at region boundaries, significantly disrupting the overall smoothness of the visualizations. Similarly, UMAP and t-SNE often struggle with low-value regions being either overshadowed by high values or averaged out, making it challenging to identify clusters with suboptimal performance. Surprisingly, Neuro-Visualizer performs the worst among the baselines for visualizing cross-modal embedding metrics. Except the Aesthetic Score, its results are highly unstable, with excessively large estimations in out-of-sample areas, and the contour maps are riddled with jagged terrains, exhibiting poor smoothness. This further underscores the strength of our AKRMap in producing more accurate and visually coherent representations. Overall, our AKRMap achieves a superior balance between local accuracy and smoothness. Furthermore, the mapping performance of traditional methods can deteriorate significantly if a smaller bandwidth is manually set, as demonstrated by additional results in Appendix C.
Figure 6: AKRMap can be used to compare generative performance of auto-regressive model and diffusion model.
Figure 7: AKRMap can be used to show the global impact of fine-tuning by comparison to base model.
5.2.2 Comparing diffusion-based model and auto-regressive model
We conduct a visual comparison of two representative T2I models from different architectural families: Stable Diffusion-v2.1 (Rombach et al., 2022) (SD-2.1), a diffusion-based model, and Infinity-2B (Han et al., 2024), an auto-regressive model. Using approximately 590,000 image captions from the MS-COCO (Lin et al., 2014) training dataset as prompts, we generate corresponding images using both models. To illustrate their performance differences, we visualize the HPSv2 score differences between Infinity and SD-2.1 (calculated as Infinity’s score minus SD-2.1’s score), as shown in Figure 6. Here, red regions highlight areas where Infinity outperforms SD-2.1 significantly, while blue regions indicate smaller differences.
The visualization reveals interesting patterns, particularly in Region A, where a significant performance gap exists between the two models. A deeper analysis reveals that this region is primarily associated with prompts related to sports players and athletes. By examining instance-level generated images with overlay feature, we find that SD-2.1 often struggles to generate human figures or produces black-and-white photographs, whereas Infinity consistently generates high-quality, colored images of athletes; see the right side of Figure 6. This indicates that Infinity demonstrates stronger capabilities in generating human figures, especially in sports-related contexts. Overall, AKRMap effectively highlights performance distribution patterns across large-scale datasets, facilitating a detailed comparative analysis of the strengths and limitations of different models.
5.2.3 Visualizing Fine-tuned Model
We showcase AKRMap’s ability to analyze the impact of model fine-tuning by comparing the SD model (Rombach et al., 2022) with Dreamlike Photoreal 2.0 (DP-2.0) (dreamlike.art, 2024), a fine-tuned variant of SD-1.5 optimized for photorealistic image generation. Our evaluation utilizes two distinct datasets: the MS-COCO (Lin et al., 2014) validation set, from which we extract 25,000 image captions as prompts, and the PartiPrompts (Yu et al., 2022) dataset, which contains 1,600 diverse prompts. Consistent with the previous analysis, we visualize the HPSv2 score differences between DP-2.0 and SD-1.5 to illustrate their comparative performance, with the results presented in Figure 7.
Overall, the effects of fine-tuning are more pronounced on the PartiPrompts benchmark, as evidenced by a larger presence of red regions in the right visualization. This is likely because MS-COCO dataset contains common captions on the web that are similar to the pretraining data of SD model, while PartiPrompts contains manually created sophisticated prompts that are challenging in different aspects. Moreover, the visualizations highlight several regions of interest (A, B, and C) where DP-2.0 demonstrates significant improvements in photorealism and overall image quality. In Region A, further analysis of individual instances reveals that DP-2.0 achieves enhanced realism in both static and dynamic objects (such as oranges and flying birds) and shows notable improvements in scenes, such as beach waves and sand textures. Region B primarily includes various stylistic renditions of raccoon images, where DP-2.0 not only accurately captures the intended styles but also excels in rendering intricate details, such as patterns on ties and textiles. Region C encompasses a variety of automotive scenes, where both the environmental contexts and the vehicles themselves exhibit greater realism and richer detail compared to SD-1.5. Importantly, our visualization demonstrates that DP-2.0 consistently outperforms SD-1.5 across the entire distribution space. This indicates that the fine-tuning process successfully enhanced the model’s photorealism while maintaining its general capabilities across other domains.
6 Discussion
Limitations. As AKRMap prioritizes metric mapping, the neighborhood preservation performance may not be comparable to traditional DR methods. Nevertheless, due to the incorporation of neighborhood constraint, AKRMap is able to maintain a desirable level of neighborhood preservation, with detailed comparison results shown in Appendix A. Furthermore, for more complex embedding metrics used in other tasks, such as those computed over sequences of embeddings (e.g., CodeBertScore (Zhou et al., 2023)), it remains uncertain how AKRMap can effectively determine a single vector representation for each instance prior to projection. This presents a potential area for further investigation.
Future Work. A recent study introduces Multi-dimensional Human Preference (MPS) (Zhang et al., 2024), which evaluates embedding scores across four dimensions: Overall, Aesthetics, Alignment, and Detail. Leveraging AKRMap to visualize multi-dimensional comparisons of different T2I models would be an interesting avenue for future exploration. AKRMap offers the potential to enhance existing cross-modal metric models through interactive human feedback. For example, we aim to fine-tune models to address domain-specific needs, such as evaluating and filtering game assets generated by T2I models.
Moreover, our method provides a versatile framework that can be adapted to other value landscape mapping challenges for high-dimensional data. For instance, AKRMap could be employed to visualize the distribution of predicted values in classical multivariate regression tasks. Beyond evaluation, AKRMap could also play a role in supporting trustworthy filtering of pretraining data by visualizing filtering scores across large-scale datasets. One pertinent example involves the use of Aesthetic Scores to filter the massive LAION-5b dataset (Schuhmann et al., 2022). We plan to scale our approach to datasets containing billions of data points, thereby increasing transparency in pretraining data selection and promoting better understanding of these processes.
7 Conclusion
In this paper, we introduce AKRMap, a dimensionality reduction method to visualize cross-modal embedding metrics through kernel regression supervised projection with adaptive generalized kernel. Based on AKRMap, we develop a visualization tool to support metric-aware visualization of cross-modal embeddings for the evaluation of text-to-image generative models. Quantitative experiments and three application scenarios show that AKRMap can facilitate trustworthy visualization of cross-modal metric for transparent evaluation.
Acknowledgments
We would like to extend our gratitude to the anonymous reviewers for their valuable comments. This work is partially supported by the National Natural Science Foundation of China (62172398, U23A20313, 62372471), The Science Foundation for Distinguished Young Scholars of Hunan Province (NO. 2023JJ10080), and Guangzhou Basic and Applied Basic Research Foundation (2024A04J6462).
Impact Statements
AKRMap advances transparent multi-modal evaluation by trustworthy visualization of cross-modal embedding metrics, which is pivotal for enhancing human-in-the-loop evaluation in tasks like T2I generation. The strategy of jointly learning discrete point DR with continuous metric landscape estimation proves highly effective in scenarios like visualizing large generated dataset, comparing different models’ performance, and revealing the global impact of fine-tuning. Our method also provides a general framework that can be adapted to embedding-based metric visualization for other tasks such as 3D generation and pretraining data filtering.
Our method highlights the synergy between automatically computed metrics and human’s interactive interpretation. With the rapid development of AI models trained and evaluated on datasets of tremendous scale, it is critical to strike a balance between the efficiency and trustworthiness of evaluation, where the transparency and visibility to humans are an essential factor.
Appendix
Appendix A Performance in Traditional Neighborhood Preservation Objective
Traditional dimensionality reduction methods often seek to preserve the neighborhood in high-dimensional space in the visualization scatterplot. Since the target of our method is not to optimize the neighborhood preservation, we do not expect our method to be better than traditional dimensionality reduction method in this regard. In Table 2, we provide statistics of traditional neighborhood trustworthiness metric (Kaski et al., 2003) to show that our method does not significantly harm the neighborhood property. Furthermore, we note that in our application where we focus on continuous metric, cluster separation is not a meaningful target since we are not looking at discrete categories as in classification tasks. As shown in Figure 8, for the HPSv2 metric on HPD dataset, even in scatterplot view AKRMap can reveal the distribution of metric significantly better than other traditional DR methods.
Table 2: Neighborhood Trustworthiness metrics for visualization of HPSv2 embeddings on HPD.
n=20 n=30 n=40 n=50
PCA 0.7243 0.7246 0.7245 0.7242
t-SNE 0.8945 0.8786 0.8667 0.8575
UMAP 0.8604 0.8467 0.8380 0.8309
AKRMap 0.8284 0.8287 0.8283 0.8277
Figure 8: Comparison of scatterplots produced by different DR methods.
Appendix B Mapping Accuracy Evaluation for in-sample points
In this section, we record the quantitative results for in-sample mapping accuracy of different methods, as shown in Table 3. The results combined with Table 1 indicate that AKRMap outperforms other methods consistently for both in-sample and out-of-sample positions.
Table 3: Quantitative comparison of mapping accuracy at in-sample points of HPD dataset.
Method CLIPScore HPSv2 PickScore Aesthetic Score
$mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$
PCA 4.0494 22.8348 5.3954 1.3744 5.3870 1.7690 1.1341 5.7472 1.4437 0.5209 10.6224 0.6749
t-SNE 4.0573 22.8162 5.3798 1.3567 5.3259 1.7606 1.1200 5.6774 1.4340 0.5192 10.5890 0.6728
UMAP 4.0906 22.8612 5.4106 1.4755 5.7803 1.9073 1.1804 5.9824 1.5145 0.5241 10.6851 0.6782
SAE 4.2591 23.2463 5.5196 1.5323 5.9966 1.9738 1.2731 6.4886 1.6270 0.5207 10.6322 0.6746
Neuro-Visualizer 10.4468 39.8135 12.1020 1.8716 7.0312 2.2237 2.0031 9.6188 2.3844 0.3102 6.1445 0.4009
AKRMap (w/o KR) 4.0506 22.6324 5.3415 1.3696 5.3706 1.7761 1.1237 5.6914 1.4343 0.3619 7.2980 0.4654
AKRMap (w/o GK) 1.6359 8.3539 2.1271 1.0479 4.0655 1.3358 0.9062 4.5448 1.1537 0.3390 6.7634 0.4320
AKRMap 1.1652 5.4238 1.5029 0.6834 2.6159 0.8980 0.6815 3.4181 0.8862 0.3142 6.1747 0.4018
Appendix C RBF Kernel Settings for Traditional DR Methods
Here we provide details and more experiment results concerning RBF kernel settings for traditional projection method. We show in Table 4 and Table 5 that traditional methods cannot achieve satisfactory performance regardless of the kernel parameter settings. Specifically, we test different commonly used bandwidth selection methods:
  • Plug-in method (Silverman) (Silverman, 2018). This is the default method we used in the paper due to its efficiency. For d-dimensional data, the Silverman’s rule is given by:
    $h=\left(\frac{4}{(d+2)n}\right)^{\frac{1}{d+4}}\hat{\sigma},$ (7)
    where $\hat{\sigma}$ is the estimated standard deviation.
  • Adaptive local bandwidth (ALB) (Cheng & Mueller, 2015). This method computes bandwidth for each point by adapting a pre-computed bandwidth based on local estimated density $f\left(x_{i}\right)$.
    $h_{i}=\lambda_{i}\times h,\ \ \lambda_{i}=(G/f(x_{i}))^{2},\ \ G=\left(\prod_{i=1}^{n}f\left(x_{i}\right)\right)^{1/n}.$ (8)
  • Cross-validation (LOOCV) (Węglarczyk, 2018). The cross-validation is performed in a Leave-one-out manner:
    $CV(h)=n^{-1}\sum_{j=1}^{n}\left[y_{j}-\hat{s}_{j}\left(x_{j}\right)\right]^{2}w\left(x_{j}\right),$ (9)
    where $\hat{s}_{j}\left(x_{j}\right)$ is the leave-one-out estimator for $y_{j}$ that is computed on all data points except $x_{j}$, and $w(x_{j})$ is a nonnegative weight function (all set to one by default). Then $h$ is selected to minimize this validation error.
Table 4: Quantitative comparison of out-of-sample mapping accuracy of different bandwidth selection methods for traditional DR.
Method CLIPScore HPSv2 PickScore
$mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$
PCA+Silverman 4.0042 16.5506 5.1142 1.4444 5.7870 1.8749 1.1497 5.9209 1.4686
PCA+ALB 4.0163 16.5445 5.1148 1.4390 5.7662 1.8651 1.1469 5.9140 1.4658
PCA+LOOCV 4.0000 16.5140 5.1034 1.4347 5.7492 1.8598 1.1415 5.8769 1.4557
t-SNE+Silverman 4.0241 16.6643 5.1523 1.4361 5.7568 1.8629 1.1579 5.9429 1.4640
t-SNE+ALB 4.0215 16.6381 5.1497 1.5147 6.1203 1.9546 1.1463 5.8838 1.4524
t-SNE+LOOCV 4.0205 16.6276 5.1366 1.5250 6.1516 1.9573 1.1520 5.9210 1.4639
UMAP+Silverman 4.0819 16.9066 5.2179 1.4432 5.7817 1.8725 1.2214 6.2721 1.5536
UMAP+ALB 4.0142 16.7284 5.1332 1.4171 5.6262 1.8285 1.2019 6.1807 1.5276
UMAP+LOOCV 4.0417 16.7337 5.1617 1.4830 5.9180 1.9233 1.2715 6.5419 1.6274
Table 5: Quantitative comparison of in-sample mapping accuracy of different bandwidth selection methods for traditional DR.
Method CLIPScore HPSv2 PickScore
$mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$ $mae$ $mape$ $rmse$
PCA+Silverman 4.0494 22.8348 5.3954 1.3744 5.3870 1.7690 1.1341 5.7472 1.4437
PCA+ALB 4.0415 22.8072 5.3806 1.3702 5.3707 1.7621 1.1284 5.7254 1.4394
PCA+LOOCV 4.0505 22.9057 5.3927 1.3747 5.3874 1.7674 1.1351 5.7502 1.4447
t-SNE+Silverman 4.0573 22.8162 5.3798 1.3567 5.3259 1.7606 1.1200 5.6774 1.4340
t-SNE+ALB 4.0476 22.7261 5.2578 1.3549 5.3168 1.7564 1.1164 5.6626 1.4292
t-SNE+LOOCV 4.0447 22.6818 5.3511 1.3545 5.3158 1.7569 1.1131 5.6390 1.4270
UMAP+Silverman 4.0906 22.8612 5.4106 1.4755 5.7803 1.9073 1.1804 5.9824 1.5145
UMAP+ALB 4.0675 22.9090 5.3831 1.4621 5.7202 1.8871 1.1729 5.9458 1.5051
UMAP+LOOCV 4.0684 22.7450 5.3680 1.4699 5.7576 1.9013 1.1730 5.9430 1.5083
In addition, we also show that it is not possible to achieve accurate mapping by manually setting smaller bandwidth than the automatically selected one. In fact, this will evidently cause severe overfitting issue and result in fragmented blocks in the map, as shown in Figure 9 of t-SNE method. This also causes the quantitative accuracy of the mapping to drop. For example, for the case in Figure 9, the error has increased as shown in Table 6.
Figure 9: Attempts to manually selecting smaller bandwidth for traditional methods may lead to worse mapping performance with severe overfitting, as exemplified by the t-SNE visualization of HPSv2 score distribution.
Table 6: Impact of manually setting smaller bandwidth.
$mae_{in}$ $mape_{in}$ $rmse_{in}$ $mae_{out}$ $mape_{out}$ $rmse_{out}$
auto-selected bandwidth 1.3567 5.3259 1.7606 1.4361 5.7568 1.8629
0.1x bandwidth 1.5403 6.0284 2.0470 1.7050 6.7994 2.1951
Appendix D Zoom in by different grid numbers
This section shows the zoom in effect of AKRMap when increasing the grid number. As shown in Figure 10, AKRMap can accurately estimate the local distribution of cross-modal metric and adds detail on demand. Specifically, with a small grid number of 30 by 30, AKRMap can already achieve the level of detail close to most baseline methods with grid number of 500 by 500. When increasing the grid number, we can see in Figure 10 that our method can accurately show the detail contour of metric distribution.
Figure 10: Zoom-in effect by different grid numbers in our method.
Appendix E Hyperparameter
In this section we discuss some hyperparameter trade-off in our method with additional experimental results.
Table 7: Experimental results on trade-off effect by $\lambda$ settings.
Method CLIPScore HPSv2 PickScore Aesthetic Score
$mae$ $rmse$ $trust$ $mae$ $rmse$ $trust$ $mae$ $rmse$ $trust$ $mae$ $rmse$ $trust$
$\lambda=0.05$ 2.0752 2.6639 0.7849 1.3613 1.7624 0.8635 1.1048 1.4027 0.8457 0.4625 0.5869 0.8731
$\lambda=0.125$ 1.8707 2.4253 0.7672 0.8108 1.1200 0.8284 0.7712 1.0225 0.8143 0.4305 0.5488 0.8704
$\lambda=0.25$ 3.4395 4.4237 0.6865 0.5621 0.8048 0.8111 0.5511 0.7413 0.7990 0.4039 0.5149 0.8626
$\lambda=0.5$ 3.4811 4.4620 0.7083 0.4453 0.6554 0.7847 0.4597 0.6346 0.7791 0.3875 0.4925 0.8545
Figure 11: AKRMap mapping results for different $\lambda$ settings.
Figure 12: Scatterplots of ablation.
Figure 13: Comparison with setting $w_{1}=0$ and $w_{2}=0$ in our MSE loss on HPSv2.
Appendix F Test on other modality embeddings
We further perform quantitative experiments on CLIP text-video embeddings on MSR-VTT dataset which consists of 200,000 text-video pairs, as well as text-audio embeddings and image-audio embeddings on Flickr 8k Audio Caption Corpus dataset consisting of 40,000 pairs.
Table 8: ImageBind image-audio embeddings mapping accuracy on Flickr 8k Audio Caption Corpus.
PCA t-SNE UMAP AKRMap
$mae$ 4.1239 3.9576 3.8811 1.9425
$mape$ 19.1321 18.3808 18.1024 8.9521
$rmse$ 5.1737 5.0254 4.8633 2.6425
Table 9: ImageBind text-audio embeddings mapping accuracy on Flickr 8k Audio Caption Corpus.
PCA t-SNE UMAP AKRMap
$mae$ 3.7194 3.5992 3.6801 1.8629
$mape$ 21.1113 20.3858 20.8498 10.1821
$rmse$ 4.7175 4.5607 4.6541 2.4841
Table 10: CLIP text-video embeddings mapping accuracy on MSR-VTT datasets
PCA t-SNE UMAP AKRMap
$mae$ 1.4908 1.4201 1.4562 0.6763
$mape$ 8.3080 7.9589 8.1636 3.4832
$rmse$ 2.1231 2.0216 2.0723 0.9578
Appendix G Weighting balancing mechanism
We also test automatic weight balancing method, using a sigmoid function to adjust the weight of KL loss.
$w(x,\mu)=\sigma(k(x-\mu))=\frac{1}{1+e^{-k(x-\mu)}},$ (10)
where $\mu$ is a threshold of acceptable $KL$ loss which we set to 2 by default and $k$ is a parameter to control the decay rate which we default to 1. Then the total loss becomes:
$L_{1}=\lambda MSE_{r}+w(KL,\mu)\cdot KL,$ (11)
where $\lambda$ is fixed to the default value 0.125.
Alternatively, for users who care more about the neighborhood preservation, we can use this method to weight the MSE term with a user-specified threshold $\mu_{1}$:
$L_{2}=w(MSE_{r},\mu_{1})\cdot\lambda MSE_{r}+\cdot KL,$ (12)
Table 11: Weight balancing results of $L_{1}$ and $L_{2}$ on HPSv2.
$mae$ $mape$ $rmse$ $trust$
w/o weight balance 0.8108 3.1935 1.1200 0.8284
$\mu=2$ 0.5118 2.0318 0.7464 0.8053
$\mu=1.8$ 0.4963 1.9632 0.7051 0.7979
$\mu=1.6$ 0.5336 2.1115 0.7772 0.8059
$\mu_{1}=1$ 1.0427 4.1393 1.3787 0.8399
$\mu_{1}=0.8$ 0.9610 3.8069 1.2686 0.8316
$\mu_{1}=0.6$ 1.0118 1.9999 1.3478 0.8426
Appendix H Interactive features evaluation
We conduct a user study and record the response time of our interactive features including zoom and overlay.
Figure 14: User study results on interactive features.